# Asymptotic Safety and Canonical Quantum Gravity

### v1

@@ 5 Canonical Quantum Gravity: from relational observables to background-independent quantization
The starting point of CQG is the Hamiltonian formulation of Einstein’s equations, typically expressed in terms of the ADM (Arnowitt-Deser-Misner) decomposition [287] which foliates spacetime into a family of spatial hypersurfaces. The canonical variables consist of the $(d-1)$-metric $q_{ij}$ on a spatial slice and its conjugate momentum related to the extrinsic curvature of the slice.
@@ 5.2 Path integral formulation
We now turn to a path integral framework. [^4] We refer the reader to Refs. 290, 291, 292, 293 for historical discussions on the subtleties relating constrained system and path integral formulation for gravitational theories. We start again from some regularized state $\omega_{\Lambda_{\text{UV}}}$, some (physical) Hamiltonian $H_{\Lambda_{\text{UV}}}$ and the related GNS data. Suppose that $H_{\Lambda_{\text{UV}}}$ is bounded from below and has a unique ground state $\Omega_{0,\Lambda_{\text{UV}}}$, i.e. $H_{\Lambda_{\text{UV}}}\Omega_{0,\Lambda_{\text{UV}}}=0$. We are interested in constructing the time ordered $n$-point correlation functions. Starting from the regularized vacuum expectation value of time-ordered Weyl operators we define:
$C_{n,\Lambda_{\text{UV}}}((t_{1},j^{1}),..,(t_{n},j^{n})):=\braket{\Omega_{0,\Lambda_{\text{UV}}}}{\mathcal{T}\left[W_{t_{n}}(j^{n},0)\cdots W_{t_{1}}(j^{1},0)\right]}{\Omega_{0,\Lambda_{\text{UV}}}}_{\mathcal{H}_{\Lambda_{\text{UV}}}}\;,$ (52)
@@
$C_{n,\Lambda_{\text{UV}}}((t_{1},a_{1}),...,(t_{n},a_{n})):=\braket{\Omega_{0,\Lambda_{\text{UV}}}}{\mathcal{T}\left[Q^{a_{n}}(t_{n})\cdots Q^{a_{1}}(t_{1})\right]}{\Omega_{0,\Lambda_{\text{UV}}}}_{\mathcal{H}_{\Lambda_{\text{UV}}}}\;.$ (53)
@@
$\chi_{\Lambda_{\text{UV}}}(j):=\braket{\Omega_{0,\Lambda_{\text{UV}}}}{\mathcal{T}\left[e^{i\int\text{d}t\,j_{a}(t)Q^{a}(t)}\right]}{\Omega_{0,\Lambda_{\text{UV}}}}_{\mathcal{H}_{\Lambda_{\text{UV}}}}\;.$ (54)
@@
[…]
@@
[…]
@@ 6.1 First examples of contact
    where the minuscule Latin letters $j=1,..,d-1$. Thus, the action depends on $2d$ scalar fields $(T,S^{j}),\;(\rho,W_{j})$. The variables are chosen to represent the physical coordinates: the $T$ scalar field serves as the time reference, it labels the proper time along the dust flow lines and effectively provides a material clock for the spacetime; the $S^{j}$ triplet of scalar fields labels the spatial coordinates along the dust congruence. The variables $\rho$ and $W_{j}$ appear as Lagrange multipliers. In particular $\rho$ is the rest-mass density of the dust, enforcing the normalization of the dust four-velocity, while the $W_{j}$ serve to enforce the condition that the dust flow lines are non-intersecting and that the spatial coordinates are transported along the dust flow. More precisely, they ensure that the spatial labels remain constant along the dust worldlines.
@@
Furthermore, by exploiting the ADM relations to connect $q_{ab}$ to $g_{\mu\nu}$ via the lapse function $N$ and the shift function $N^{a}$ in (4) one can change variables in the functional integration in the path integral in order to obtain an integration over $[dg]\mathcal{M}^{\prime}[g]$, where the measure gets now accordingly modified. This comes with the advantage that we can use standard covariant heat kernel methods as in (41) to compute the traces in the flow equation (17).
@@
In this second investigation, the authors accounted for the non-trivial measure by implementing an appropriate field redefinition (see Ref. 300 for a recent discussion on the meaning of field redefinitions)
@@
Importantly a new versatile tool was used, imported from loops computations in Feynman integrals: the Mellin-Barnes integrals [308, 309, 310, 311]. This allowed a controlled evaluation of the higher order contributions in the proper time integration (see Appendix B in Ref. 286).
@@ 6.2 Novel features from asymptotically safe canonical quantum gravity
Moreover, the flow does not present the typical spiraling and complex conjugated critical exponents as those presented in Section 4. Finally, due to numerical approximations needed to solve the flow detailed in Ref. 281, the trajectories do not reach the Gaussian fixed point. However, this result arises solely as an artifact of numerical approximation and does not reflect a fundamental property of the system. On the other side, there is no singular region for positive $\lambda>1/2$, meaning that those trajectories going towards $\lambda\to\infty$ for decreasing $k$ can be integrated down to $k\to 0$, reaching the effective regime.
@@
Finally, this investigation opens new intriguing questions regarding the nature of the RG. It appears that the relational nature of the underlying theory translates into a relational RG. As we mentioned in the section before, a given universality is associated with critical phenomena. However, in the framework discussed here, where the gauge degrees of freedom are discarded from the beginning, one can question whether different reference frames would correspond to the same universality class. Although this is something expected once the proper transformation between reference frames has been taken into account, it might be challenging to realize, especially in a reduced phase space. Moreover, in certain situations where the reference frames are “non-ideal” [324] and/or the causal structure of the underlying quantum spacetime exhibits specific features, some frames may be better suited than others for detecting the observables (see Refs. 325, 326, 327 for some examples in a perturbative cosmological setting). On a speculative level, this could imply that, in such “non-ideal” material reference frames, the theory may fail to exhibit a fixed point. These aspects open up promising directions for future research and will be explored in more detail in forthcoming work.
@@ 7 Conclusions and Outlook
We have highlighted how certain aspects of the path integral must be adapted in order to account for a canonical theory of quantum gravity, by relaxing explicit covariance and defining the path integral in a relational manner. The path integral and in particular the effective action, was then constructed by means of the FRG setup. The flow equation is used as a tool to reach the $k\to 0$ limit, where all the quantum fluctuations have been integrated out.
It is expected that this relational, non-covariant path integral can be connected to different relational, non-covariant path integrals through suitable transformations that account for changes of the observer. This is a mathematically challenging task and should be explored in future work. Such non-covariance is not necessarily troubling since it arises from the physical characteristics of observables and observers in GR and as long as transformations between reference systems remain well-defined. This can potentially be achieved at the level of the unreduced phase space by leveraging the framework developed in the context of quantum reference frames [328, 329, 330]. On the other hand, we have outlined in Section 3 how initial steps have been made toward defining a covariant theory from Hamiltonian quantization. This approach could serve as a foundation for future work, aiming to bridge the two frameworks and apply standard AS techniques to the resulting path integral.
@@ Acknowledgments
I am grateful to Thomas Thiemann for fruitful discussions and to Alfio Bonanno, Muxin Han, Roberto Percacci and especially Frank Saueressig for fruitful discussions and comments on the manuscript. I am supported by the Friedrich-Alexander University (FAU) Emerging Talents Initiative (ETI).

### v2

@@ 5 Canonical Quantum Gravity: from relational observables to background-independent quantization
The starting point of CQG is the Hamiltonian formulation of Einstein’s equations, typically expressed in terms of the ADM (Arnowitt-Deser-Misner) decomposition [317] which foliates spacetime into a family of spatial hypersurfaces. The canonical variables consist of the (d1)(d-1)-metric $q_{ij}$ on a spatial slice and its conjugate momentum related to the extrinsic curvature of the slice.
@@ 5.2 Path integral formulation
We now turn to a path integral framework. [^4] We refer the reader to Refs. 320, 321, 322, 323 for historical discussions on the subtleties relating constrained system and path integral formulation for gravitational theories. We start again from some regularized state $\omega_{\Lambda_{\text{UV}}}$, some (physical) Hamiltonian $H_{\Lambda_{\text{UV}}}$ and the related GNS data. Suppose that $H_{\Lambda_{\text{UV}}}$ is bounded from below and has a unique ground state $\Omega_{0,\Lambda_{\text{UV}}}$, i.e. $H_{\Lambda_{\text{UV}}}\Omega_{0,\Lambda_{\text{UV}}}=0$. We are interested in constructing the time ordered $n$-point correlation functions. Starting from the regularized vacuum expectation value of time-ordered Weyl operators we define:
$C_{n,\Lambda_{\text{UV}}}((t_{1},j^{1}),..,(t_{n},j^{n})):=\braket{\Omega_{0,\Lambda_{\text{UV}}}|\mathcal{T}\left[W_{t_{n}}(j^{n},0)\cdots W_{t_{1}}(j^{1},0)\right]|\Omega_{0,\Lambda_{\text{UV}}}}_{\mathcal{H}_{\Lambda_{\text{UV}}}}\;,$ (52)
@@
$C_{n,\Lambda_{\text{UV}}}((t_{1},a_{1}),...,(t_{n},a_{n})):=\braket{\Omega_{0,\Lambda_{\text{UV}}}|\mathcal{T}\left[Q^{a_{n}}(t_{n})\cdots Q^{a_{1}}(t_{1})\right]|\Omega_{0,\Lambda_{\text{UV}}}}_{\mathcal{H}_{\Lambda_{\text{UV}}}}\;.$ (53)
@@
$\chi_{\Lambda_{\text{UV}}}(j):=\braket{\Omega_{0,\Lambda_{\text{UV}}}|\mathcal{T}\left[e^{i\int\text{d}t\,j_{a}(t)Q^{a}(t)}\right]|\Omega_{0,\Lambda_{\text{UV}}}}_{\mathcal{H}_{\Lambda_{\text{UV}}}}\;.$ (54)
@@
$\braket{\Omega_{0,\Lambda_{\text{UV}}}|\;\cdot\;|\Omega_{0,\Lambda_{\text{UV}}}}=\lim_{T\to\infty}\frac{\braket{\Omega_{\Lambda_{\text{UV}}}|U(T)\;\cdot\;U(-T)|\Omega_{\Lambda_{\text{UV}}}}}{\braket{\Omega_{\Lambda_{\text{UV}}}|U(-2T)|\Omega_{\Lambda_{\text{UV}}}}}\;.$ (55)
@@
$Z(j)=\int\;[dq]\;\mathcal{M}[q]\;\overline{\Omega(q(\infty))}\;\Omega(q(-\infty))\;e^{i\int\;dt\;j_{a}(t)q^{a}(t)}\;e^{-\int\;dtL(q(t),\dot{q}(t))}\;,$ (58)
@@ 6.1 First examples of contact
    where the minuscule Latin letters j=1,..,d1j=1,..,d-1. Thus, the action depends on $2d$ scalar fields $(T,S^{j}),\;(\rho,W_{j})$. The variables are chosen to represent the physical coordinates: the $T$ scalar field serves as the time reference, it labels the proper time along the dust flow lines and effectively provides a material clock for the spacetime; the $S^{j}$ triplet of scalar fields labels the spatial coordinates along the dust congruence. The variables $\rho$ and $W_{j}$ appear as Lagrange multipliers. In particular $\rho$ is the rest-mass density of the dust, enforcing the normalization of the dust four-velocity, while the $W_{j}$ serve to enforce the condition that the dust flow lines are non-intersecting and that the spatial coordinates are transported along the dust flow. More precisely, they ensure that the spatial labels remain constant along the dust worldlines.
@@
Furthermore, by exploiting the ADM relations to connect $q_{ab}$ to $g_{\mu\nu}$ via the lapse function $N$ and the shift function $N^{a}$ in (4) one can change variables in the functional integration in the path integral in order to obtain an integration over $dg]\mathcal{M}^{\prime}[g]$, where the measure gets now accordingly modified. This comes with the advantage that we can use standard covariant heat kernel methods as in (41) to compute the traces in the flow equation (17).
@@
In this second investigation, the authors accounted for the non-trivial measure by implementing an appropriate field redefinition (see Ref. 330 for a recent discussion on the meaning of field redefinitions)
@@
Importantly a new versatile tool was used, imported from loops computations in Feynman integrals: the Mellin-Barnes integrals [338, 339, 340, 341]. This allowed a controlled evaluation of the higher order contributions in the proper time integration (see Appendix B in Ref. 316).
@@ 6.2 Novel features from asymptotically safe canonical quantum gravity
Moreover, the flow does not present the typical spiraling and complex conjugated critical exponents as those presented in Section 4. Finally, due to numerical approximations needed to solve the flow detailed in Ref. 311, the trajectories do not reach the Gaussian fixed point. However, this result arises solely as an artifact of numerical approximation and does not reflect a fundamental property of the system. On the other side, there is no singular region for positive $\lambda>1/2$, meaning that those trajectories going towards $\lambda\to\infty$ for decreasing $k$ can be integrated down to $k\to 0$, reaching the effective regime.
@@
Finally, this investigation opens new intriguing questions regarding the nature of the RG. It appears that the relational nature of the underlying theory translates into a relational RG. As we mentioned in the section before, a given universality is associated with critical phenomena. Indeed, for all we know, quantities derived via a relational strategy in a particular physical frame will generally depend on that frame, and it is not yet fully understood what aspects of the dynamics, such as the RG flow, remain invariant under a change of physical frame, or what class of transformations relates observables defined in different frames. These transformations are not diffeomorphisms by construction, since the reference fields themselves carry physical meaning. Clarifying their nature and the possible equivalence of frames is therefore an important open issue, though not one that implies any tension with covariance or diffeomorphism invariance. In approximate regimes involving idealized reference systems, neglecting backreaction and quantum properties, the physical frame effectively corresponds to a specific gauge or coordinate choice, and in such cases the expected gauge invariance and frame equivalence of results are recovered.
In particular, in the reduced framework discussed here, where the gauge degrees of freedom are discarded from the beginning, one can question whether different reference frames would correspond to the same universality class. Although this is something expected once the proper transformation between reference frames has been taken into account, it might be challenging to realize, especially in a reduced phase space. Moreover, in certain situations where the reference frames are “non-ideal” [354] and/or the causal structure of the underlying quantum spacetime exhibits specific features, some frames may be better suited than others for detecting the observables (see Refs. 355, 356, 357 for some examples in a perturbative cosmological setting). On a speculative level, this could imply that, in such “non-ideal” material reference frames, the theory may fail to exhibit a fixed point. These aspects open up promising directions for future research and will be explored in more detail in forthcoming work.
@@ 7 Conclusions and Outlook
We have highlighted how certain aspects of the path integral can be adapted in order to account for a canonical theory of quantum gravity, by relaxing explicit covariance and defining the path integral in a relational manner. The path integral and in particular the effective action, was then constructed by means of the FRG setup. The flow equation is used as a tool to reach the $k\to 0$ limit, where all the quantum fluctuations have been integrated out.
It is expected that this relational, non-covariant path integral can be connected to different relational, non-covariant path integrals through suitable transformations that account for changes of the observer. This is a mathematically challenging task and should be explored in future work. Such non-covariance is not necessarily troubling since it arises from the physical characteristics of observables and observers in GR and as long as transformations between reference systems remain well-defined. This can potentially be achieved at the level of the unreduced phase space by leveraging the framework developed in the context of quantum reference frames [358, 359, 360] and TGFT cosmology [243, 148]. On the other hand, we have outlined in Section 3 how initial steps have been made toward defining a covariant theory from Hamiltonian quantization. This approach could serve as a foundation for future work, aiming to bridge the two frameworks and apply standard AS techniques to the resulting path integral.
@@ Acknowledgments
I am grateful to Thomas Thiemann for fruitful discussions and to Alfio Bonanno, Muxin Han, Roberto Percacci and especially Frank Saueressig for fruitful discussions and comments on the manuscript. I thank Daniele Oriti for valuable comments that clarified the link between canonical and covariant quantum gravity models and the relational strategy’s compatibility with general covariance. I am supported by the Friedrich-Alexander University (FAU) Emerging Talents Initiative (ETI).